OpenStudy (anonymous):

find the x-values that are solutions of the equation 5cot^2x-15=0

OpenStudy (anonymous):

5cot^2x-15=0 cot^2x=15/ 5=3 => cotx = sqrt(3)

OpenStudy (anonymous):